OpenStudy (anonymous):

I need help checking this I c if possible can you tell me where i went wrong

OpenStudy (anonymous):

hold on i have to post it

OpenStudy (anonymous):

here is what i got Area=lw= (t+5)(t+3) = t^2+8t+15 Area=t^(2+) 5t+3t=15 = t^2+8t+15

OpenStudy (anonymous):

here is the question Find two algebraic expressions for the area of each figure. First, regard the figure as one large rectangle, and then regard the figure as a sum of four smaller rectangles.

OpenStudy (anonymous):